A simple random sample of 60 items resulted in a sample mean of 90. The population standard deviation is 13.

1)  Compute the 95% confidence interval for the population mean (to 1 decimal).

2) Assume that the same sample mean was obtained from a sample of 120 items. Provide a 95% confidence interval for the population mean (to 2 decimals).

3) What is the effect of a larger sample size on the margin of error?
Answer #1

Given

n=60

Sample mean \bar{x} = 90

population standard deviation \sigma = 13

1)

Compute the 95% confidence interval for the population mean

z value at 0.05 = 1.96

=\bar{x}\pm 1.96*\frac{\sigma }{\sqrt{n}}

=90\pm 1.96*\frac{13}{\sqrt{60}}

=90\pm 3.96

=86.71,93.96

2)

Here

n = 120

=\bar{x}\pm 1.96*\frac{\sigma }{\sqrt{n}}

=90\pm 1.96*\frac{13}{\sqrt{120}}

=90\pm 2.326

=87.674,92.326

3)

The margin of error decreases when the Sample size increases.
Answer #1

Given

n=60

Sample mean \bar{x} = 90

population standard deviation \sigma = 13

1)

Compute the 95% confidence interval for the population mean

z value at 0.05 = 1.96

=\bar{x}\pm 1.96*\frac{\sigma }{\sqrt{n}}

=90\pm 1.96*\frac{13}{\sqrt{60}}

=90\pm 3.96

=86.71,93.96

2)

Here

n = 120

=\bar{x}\pm 1.96*\frac{\sigma }{\sqrt{n}}

=90\pm 1.96*\frac{13}{\sqrt{120}}

=90\pm 2.326

=87.674,92.326

3)

The margin of error decreases when the Sample size increases.